# High-low analysis and small cap decoupling over non-Archimedean local fields

### v1

@@ 2.4 Reduction to local estimates
$(\mathfrak{s}^{k_{*}})^{\frac{p}{2}-\frac{p}{q}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-\frac{p}{q}-3}\leq R^{\beta(\frac{p}{2}-\frac{p}{q})}.$
@@
$(\mathfrak{s}^{k_{*}})^{\frac{p}{2}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-3}\leq R^{\beta(\frac{p}{2}-\frac{p}{q})}(\#\gamma)^{\frac{p}{q}},$
@@
$\begin{split}\mathfrak{a}^{p}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}&\leq C2^{\frac{p}{2}-3}e^{2p-4}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(\frac{p}{2}-\frac{p}{q})}(\#\gamma)^{\frac{p}{q}}\lambda^{p}\mathfrak{r}\cdot R^{1+\beta}\\ &\leq C2^{\frac{p}{2}-3}e^{4p-3}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(\frac{p}{2}-\frac{p}{q})}\Big{(}\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.\end{split}$
Proposition 3.5.
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 8e^{4}\mathbf{p}^{5}N^{8}R^{\beta}.$
Proof.
By repeated applications of the narrow lemma,
$|f(x)|\leq\left(1+\frac{1}{N-1}\right)^{N}\max_{\theta}|f_{\theta}(x)|+\mathbf{p}N\sum_{k=1}^{N}\left(1+\frac{1}{N-1}\right)^{k-1}\max_{\tau_{k-1}}\max_{\tau_{k}\neq\tau_{k}^{\prime}\prec\tau_{k-1}}|f_{\tau_{k}}(x)f_{\tau_{k}^{\prime}}(x)|^{1/2},$
so that
$|f(x)|^{4}\leq 2^{3}\left(1+\frac{1}{N-1}\right)^{4N}\max_{\theta}|f_{\theta}(x)|^{4}+\mathbf{p}^{4}N^{4}\sum_{k=1}^{N}2^{3}N^{3}\left(1+\frac{1}{N-1}\right)^{4(k-1)}\max_{\tau_{k-1}}\max_{\tau_{k}\neq\tau_{k}^{\prime}\prec\tau_{k-1}}|f_{\tau_{k}}(x)f_{\tau_{k}^{\prime}}(x)|^{2}.$
For each $\theta$,
$\|f_{\theta}\|_{4}^{4}\leq R^{2(\beta-\frac{1}{2})}\sum_{\gamma\prec\theta}\|f_{\gamma}\|_{4}^{4}.$
Choose now some $\tau_{k-1}$. Let $(c,c^{2})\in\tau_{k-1}$, and write
$g(x,y)=f_{\tau_{k-1}}(\varpi^{-\eta(k-1)}(x-c),\varpi^{-2\eta(k-1)}(y-c^{2})),$
$g_{1}(x,y)=f_{\tau_{k}}(\varpi^{-\eta(k-1)}(x-c),\varpi^{-2\eta(k-1)}(y-c^{2})),\quad g_{2}(x,y)=f_{\tau_{k}^{\prime}}(\varpi^{-\eta(k-1)}(x-c),\varpi^{-2\eta(k-1)}(y-c^{2}))$
Then $g$ has spectral support in $\mathcal{N}_{R_{k-1}^{2}R^{-1}}(\mathbb{P}^{1})$, so by Córdoba-Fefferman we have
$\int|g_{1}g_{2}|^{2}\leq 2R^{\beta}R_{k-1}^{-2\beta}\sum_{\tilde{\theta}}\|g_{\tilde{\theta}}\|_{4}^{4},$
where the $\tilde{\theta}$ have dimensions $R^{-\beta}R_{k-1}^{\beta}\times R^{-1}R_{k-1}$. By flat decoupling,
$\|g_{\tilde{\theta}}\|_{4}^{4}\leq R_{k-1}^{2(2\beta-1)}\sum_{\tilde{\gamma}}\|g_{\tilde{\gamma}}\|_{4}^{4}$
Thus,
$\int|f_{\tau_{k}}f_{\tau_{k}^{\prime}}|^{2}\leq 2R^{\beta}R_{k-1}^{2\beta-2}\sum_{\gamma\prec\tau_{k}\sqcup\tau_{k}^{\prime}}\int|f_{\gamma}|^{4}.$
We have reached the estimate
$\int|f|^{4}\leq 2^{3}\left(1+\frac{1}{N-1}\right)^{4N}R^{2\beta-1}\sum_{\gamma}\|f_{\gamma}\|_{4}^{4}+16\mathbf{p}^{5}N^{7}R^{\beta}\sum_{k=1}^{N}\left(1+\frac{1}{N-1}\right)^{4(k-1)}R_{k-1}^{2\beta-2}\sum_{\gamma}\|f_{\gamma}\|_{4}^{4},$
i.e.
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 8e^{4}\mathbf{p}^{5}N^{8}R^{\beta}.$

### v2

@@ 2.4 Reduction to local estimates
$(\mathfrak{s}^{k_{*}})^{\frac{r}{2}-\frac{r}{q}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-\frac{r}{q}-3}\leq R^{\beta(\frac{r}{2}-\frac{r}{q})}.$
@@
$(\mathfrak{s}^{k_{*}})^{\frac{r}{2}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-3}\leq R^{\beta(\frac{r}{2}-\frac{r}{q})}(\#\gamma)^{\frac{r}{q}},$
@@
$\begin{split}\mathfrak{a}^{r}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)&\leq C2^{\frac{r}{2}-3}e^{2r-4}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(\frac{r}{2}-\frac{r}{q})}(\#\gamma)^{\frac{r}{q}}\lambda^{r}\mathfrak{t}\cdot R^{1+\beta}\\ &\leq C2^{\frac{r}{2}-3}e^{4r-3}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(\frac{r}{2}-\frac{r}{q})}\Big(\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{q}\Big)^{\frac{r}{q}}.\end{split}$
4 Transference and real integrals
In this section, we supply the tools needed to convert our main results (Theorems (1.1) and (1.2)) into the applications described in subsection (1.1). The basic procedure was already described in our previous work [13]; we give a simplified account here to be self-contained.
We begin by sketching a proof of Theorem (1.7). It will contain many of the details needed to explain the other two applications.
Proof sketch of Theorem (1.7).
We will first handle the case $\gamma=0$, and then explain the needed modifications for the general $\gamma$ case.
Let $N\in\mathbb{N}$ and $\sigma\in[0,1]$ be such that $N,N^{\sigma}\in p^{\mathbb{Z}}$ with $p$ an odd prime. Fix $\mathbb{K}=\mathbb{Q}_{p}$. Consider the domain $\mathfrak{D}\subseteq\mathbb{Q}_{p}^{2}$ defined by
$\mathfrak{D}=\mathcal{O}\times B(0,N^{\sigma-1}).$
Thus, $\mathfrak{D}$ consists of those $(x,y)\in\mathbb{Q}_{p}^{2}$ with $|x|\leq 1$, $|y|\leq N^{\sigma-1}$. By appealing to Theorem (1.1), we may show
$\fint_{\mathfrak{D}}\Big|\sum_{n=0}^{N-1}a_{n}e_{\mathbb{Q}_{p}}\Big(\frac{n}{N}x+\frac{n^{2}}{N^{2}}t\Big)\Big|^{r}dxdt\lesssim_{\mathbf{p}}(\log N)^{O(1)}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\left(\sum_{n=0}^{N-1}|a_{n}|^{q}\right)^{r/q}.$ (4.1)
Here we have written $e=e_{\mathbb{Q}_{p}}$ for the function specified in (1.10). The bound (4.1) follows from standard changes-of-variable and periodicity, together with an application of small cap decoupling (1.1).
Next, note that
$f(x,t)=\sum_{n=0}^{N-1}a_{n}e_{\mathbb{Q}_{p}}\Big(\frac{n}{N}x+\frac{n^{2}}{N^{2}}t\Big)$
has the local constancy property $f(x+h,t+k)=f(x,t)$ whenever $|h|\leq N^{-1},|t|\leq N^{-2}$. In particular, our integral may be written as the following Riemann sum:
$\begin{split}\fint_{\mathfrak{D}}|f(x,t)|^{r}dxdt&=N^{-2-\sigma}\sum_{a=0}^{N-1}\sum_{b=0}^{N^{1+\sigma}-1}|f(a,N^{1-\sigma}b)|^{r}\\ &=N^{-2-\sigma}\sum_{a=0}^{N-1}\sum_{b=0}^{N^{1+\sigma}-1}\Bigg|\sum_{n=0}^{N-1}a_{n}e_{\mathbb{Q}_{p}}\Big(\frac{na}{N}+\frac{n^{2}b}{N^{1+\sigma}}\Big)\Bigg|^{r}.\end{split}$
Each of the arguments in the latter $e_{\mathbb{Q}_{p}}$ is a rational number whose denominator is a power of $p$. Thus, after inspecting (1.10), we may freely substitute $e_{\mathbb{Q}_{p}}$ for $e_{\mathbb{R}}$:
$\fint_{\mathfrak{D}}|f(x,t)|^{r}dxdt=N^{-2-\sigma}\sum_{a=0}^{N-1}\sum_{b=0}^{N^{1+\sigma}-1}\Bigg|\sum_{n=0}^{N-1}a_{n}e_{\mathbb{R}}\Big(\frac{na}{N}+\frac{n^{2}b}{N^{1+\sigma}}\Big)\Bigg|^{r}.$
Here of course $e_{\mathbb{R}}$ is the usual Euclidean phase $e_{\mathbb{R}}(x)=\exp(2\pi ix)$. In particular, we may write
$\fint_{\mathfrak{D}}|f(x,t)|^{r}dxdt=\int\Bigg|\sum_{n=0}^{N-1}a_{n}e_{\mathbb{R}}\Big(nx+n^{2}t\Big)\Bigg|^{r}d\Delta(x,t),$
where we have written $\Delta$ for the measure on $\mathbb{R}^{2}$ defined by
$\Delta=N^{-2-\sigma}\sum_{a=0}^{N-1}\sum_{b=0}^{N^{1+\sigma}-1}\delta_{\frac{a}{N}}\otimes\delta_{\frac{b}{N^{1+\sigma}}}.$
Here $\delta_{c}$ is the Dirac delta mass located at $c\in\mathbb{R}$. Since translations $(x,t)\mapsto(x+\mathfrak{x},t+\mathfrak{t})$ amount to replacing the coefficients $a_{n}$ with unimodular multiples $\tilde{a}_{n}$, $|a_{n}|=|\tilde{a}_{n}|$, we may write
$\fint_{A(N,\sigma;0)}\Bigg|\sum_{n=1}^{N}a_{n}e\Big(nx+n^{2}t\Big)\Bigg|^{r}dxdt\leq\max_{\begin{subarray}{c}\{\tilde{a}_{n}\}_{n}\\ |\tilde{a}_{n}|=|a_{n}|\end{subarray}}\int\Bigg|\sum_{n=0}^{N-1}\tilde{a}_{n}e\Big(nx+n^{2}t\Big)\Bigg|^{r}d\Delta(x,t).$
In particular, we obtain the estimate
$\fint_{A(N,\sigma;0)}\Bigg|\sum_{n=1}^{N}a_{n}e\Big(nx+n^{2}t\Big)\Bigg|^{r}dxdt\lesssim_{\mathbf{p}}(\log N)^{O(1)}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\left(\sum_{n=0}^{N-1}|a_{n}|^{q}\right)^{r/q}.$
We have justified Theorem (1.7) when $\gamma=0$. We now discuss the intermediate case $\gamma\in(0,1)$; recall that the $\gamma=1$ case is the usual “short mean value estimate” à la Theorem (1.3), and the logarithmic factor for (real) small cap decoupling was obtained in [12]. The basic point is to consider exponential sums of the form
$f(x,t,x^{\prime},t^{\prime})=\sum_{n=1}^{N}a_{n}e_{\mathbb{Q}_{p}}(nx+n^{2}t)e_{\mathbb{R}}\Big(\frac{n}{N}x^{\prime}+\frac{n^{2}}{N^{2}}t^{\prime}\Big),$
where $x,t\in\mathbb{Q}_{p}$ and $x^{\prime},t^{\prime}\in\mathbb{R}$. Thus, $f$ is a complex-valued function defined on the LCA group $\mathbb{Q}_{p}\times\mathbb{R}$; the general theory of Pontryagin duality implies that the tools of Fourier analysis on the $\mathbb{Q}_{p}$ and $\mathbb{R}$ factors may be freely lifted to the product.
We supply several of the technical steps. Let
$g(x,t,x^{\prime},t^{\prime})=\sum_{n=1}^{N}a_{n}1_{N^{-1}\mathbb{Z}_{p}^{2}}(x,t)\eta_{B_{N}}(x^{\prime},t^{\prime})e_{\mathbb{Q}_{p}}(nx+n^{2}t)e_{\mathbb{R}}\Big(\frac{n}{N}x^{\prime}+\frac{n^{2}}{N^{2}}t^{\prime}\Big),$
where $\eta_{B_{N}}$ is a nonnegative smooth function satisfying $\eta_{B_{N}}\gtrsim 1$ on $B_{N}\subseteq\mathbb{R}^{2}$ and $\hat{\eta}_{B_{N}}$ is supported in $B_{N^{-1}}$. Let $g_{n}$ be the $n$th summand in the definition of $g$. By using $p$-adic decoupling, Minkowski, Fubini, and real decoupling in turn,
$\begin{split}\int_{\mathbb{R}^{2}}&\int_{\mathbb{Q}_{p}^{2}}|g(x,t,x^{\prime},t^{\prime})|^{r}dxdtdx^{\prime}dt^{\prime}\\ &\leq p^{7}B(N^{1-\gamma})\int_{\mathbb{R}^{2}}\left(\sum_{\tau}\left(\int_{\mathbb{Q}_{p}^{2}}\Bigg|\sum_{n\in\tau}g_{n}(x,t,x^{\prime},t^{\prime})\Bigg|^{r}dxdt\right)^{q/r}\right)^{r/q}dx^{\prime}dt^{\prime}\\ &\leq p^{7}B(N^{1-\gamma})\left(\sum_{\tau}\left(\int_{\mathbb{Q}_{p}^{2}}\int_{\mathbb{R}^{2}}\Bigg|\sum_{n\in\tau}g_{n}(x,t,x^{\prime},t^{\prime})\Bigg|^{r}dx^{\prime}dt^{\prime}dxdt\right)^{q/r}\right)^{r/q}\\ &\leq p^{7}B(N^{1-\gamma})B(N^{\gamma})\left(\sum_{\tau}\sum_{\tau^{\prime}}\left(\int_{\mathbb{Q}_{p}^{2}}\int_{\mathbb{R}^{2}}\Bigg|\sum_{n\in\tau\cap\tau^{\prime}}g_{n}(x,t,x^{\prime},t^{\prime})\Bigg|^{r}dx^{\prime}dt^{\prime}dxdt\right)^{q/r}\right)^{r/q}.\end{split}$
Here we have the notation
$B(M)=C(\log M)^{C^{\prime}}\big(M^{(r-\frac{r}{q}-1)-\sigma-1}+M^{\frac{r}{2}-\frac{r}{q}}\big),$
with $C,C^{\prime}\gg 1$ constants arising from the decoupling theorem (1.1) and the real analogue in [12]. The sets $\tau$ range over $\mathcal{P}(\mathcal{O},N^{\gamma-1})$ and the sets $\tau^{\prime}$ range over the partition of $[0,N]$ into intervals of length $N^{1-\gamma}$. It follows that the intersections $\tau\cap\tau^{\prime}$ are singletons, so
$\int_{\mathbb{R}^{2}}\int_{\mathbb{Q}_{p}^{2}}|g(x,t,x^{\prime},t^{\prime})|^{r}dxdtdx^{\prime}dt^{\prime}\leq p^{7}B(N)N^{4}\left(\sum_{n=1}^{N}|a_{n}|^{q}\right)^{r/q}.$
Applying the same rescaling, transference, and translation principles as before, we obtain the estimate
$\fint_{A(N,\sigma;\gamma)}\Bigg|\sum_{n=1}^{N}e(nx+n^{2}t)\Bigg|^{r}dxdt\lesssim p^{7}(\log N)^{O(1)}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\left(\sum_{n=1}^{N}|a_{n}|^{q}\right)^{r/q},$
as desired.
We sketch the argument for Theorem (1.10) next, which is similar in spirit to the previous result.
Proof of Theorem (1.10).
We work over $\mathbb{K}=\mathbb{Q}_{3}\big[\sqrt{-1}\big]$. A uniformizer $\omega$ for $\mathbb{K}$ is $3$; here, we have used that $p\equiv 3$ mod $4$. Our choice of basic character is
$e_{\mathbb{K}}(x)=e_{\mathbb{Q}_{3}}\big(\mathrm{Tr}_{\mathbb{K}/\mathbb{Q}_{3}}(x)\big).$
This $e_{\mathbb{K}}$ satisfies the necessary condition $e_{\mathbb{K}}(\mathcal{O})=\{1\}\neq e_{\mathbb{K}}(3^{-1}\mathcal{O})$. We also have
$e_{\mathbb{K}}\big(a+b\sqrt{-1}\big)=e_{\mathbb{Q}_{3}}(a),\quad a,b\in\mathbb{Q}_{3}.$
We study the same mean-value estimate considered in the proof of Theorem (1.7), framed over $\mathbb{K}$ instead of $\mathbb{Q}_{p}$. We have
$\begin{split}\fint_{\mathcal{O}\times B(0,N^{\sigma-1})}&\Bigg|\sum_{n,m=1}^{N}a_{n,m}e_{\mathbb{K}}\Big(\frac{n+m\sqrt{-1}}{N}x+\frac{(n+m\sqrt{-1})^{2}}{N^{2}}t\Big)\Bigg|^{r}\\ &\lesssim(\log N)^{O(1)}\big(N^{(r-\frac{r}{q}-1)-1-\sigma}+N^{\frac{r}{2}-\frac{r}{q}}\big)\left(\sum_{n,m=1}^{N}|a_{n,m}|^{q}\right)^{r/q}\end{split}$
As in the previous transference argument, we may reduce the integral to a Riemann sum. Here, we select $x$ to run over $a+b\sqrt{-1}$, $0\leq a,b\leq N-1$, and $t$ to run over $c+d\sqrt{-1}$, $0\leq c,d\leq N^{1+\sigma}-1$. Using the choice of $e_{\mathbb{K}}$, we see that
$e_{\mathbb{K}}\Big(\frac{n+m\sqrt{-1}}{N}x+\frac{(n+m\sqrt{-1})^{2}}{N^{2}}t\Big)=e_{\mathbb{R}}\Big(\frac{na}{N}-\frac{mb}{N}+\frac{(n^{2}-m^{2})c}{N^{2}}-\frac{2nmd}{N^{2}}\Big).$
The remainder of the transference goes through without substantial modification. ∎
We end by giving a proof of Theorem (1.9).
Proof of Theorem (1.9).
Define the function
$g(x,t)=1_{N^{-1-\sigma}\mathbb{Z}_{p}^{2}}(x,t)\sum_{j\in S}a_{j}e_{\mathbb{Q}_{p}}\Big(jx+j^{2}t\Big).$
Here $\mathbb{Q}_{p}$ is determined by the choice of $p$ in the statement of the theorem. Thus, $g$ is a Schwartz–Bruhat function on $\mathbb{Q}_{p}^{2}$.
By Theorem (1.2), we have
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\})\leq\frac{1}{50}p^{7}(\log_{p}N)^{10}\sum_{\begin{subarray}{c}k\in p^{\mathbb{Z}}\\ N^{-\frac{1+\sigma}{2}}\leq p^{-k}\leq 1\end{subarray}}\sum_{\begin{subarray}{c}\tau\\ \mathrm{diam}(\tau)=p^{-k}\end{subarray}}\sum_{U\in\mathcal{G}_{\tau}[g;\alpha]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau}|g_{\theta}|^{2}\right)^{2},$
where the $\theta$ are the caps of diameter $N^{-\frac{1+\sigma}{2}}$. By Proposition (3.1), for each such $k$ and $\tau$ we have
$\sum_{U\in\mathcal{G}_{\tau}[g;\alpha]}\mu(U)\left(\fint_{U}\sum_{\theta\prec\tau}|g_{\theta}|^{2}\right)^{2}\leq\max_{\gamma_{k}\prec\tau}\#(S\cap\gamma_{k})\#(S\cap\tau)\sum_{j\in S\cap\tau}N^{2+2\sigma}|a_{j}|^{4},$
where $\gamma_{k}$ are caps of diameter $p^{k}N^{-1-\sigma}$. Summing over $\tau$, we obtain
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\})\leq\frac{1}{100}p^{7}(\log_{p}N)^{11}\alpha^{-4}N^{2+2\sigma}\max_{k}\Big(\lambda_{k}(S)\cdot\lambda_{(1+\sigma)\log_{p}N-k}(S)\Big)\sum_{j\in S}|a_{j}|^{4}.$ (4.2)
It remains to relate $\mu(\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\})$ to $\mathcal{L}^{2}\big(U_{\alpha}\cap A(N,\sigma;0)\big)$. If $0\leq a\leq N-1$ and $0\leq b\leq N^{1+\sigma}-1$ are such that $(\frac{a}{N},\frac{b}{N^{1+\sigma}})\in\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\}$, then
$\left|\sum_{j\in S}a_{j}e_{\mathbb{R}}\Big(j\frac{a}{N}+j^{2}\frac{b}{N^{1+\sigma}}\Big)\right|>\alpha.$
Let
$\mathcal{A}=\Big\{0\leq a\leq N-1,0\leq b\leq N^{1+\sigma}-1:\big(\frac{a}{N},\frac{b}{N^{1+\sigma}}\big)\in\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\}\Big\}.$
Then, by local constancy and periodicity, we find that
$\mu(\{x\in\mathbb{Q}_{p}^{2}:|g(x)|>\alpha\})=N^{\sigma}\#\mathcal{A}.$
To get back to $U_{\alpha}$, we will need to study random perturbations of our main estimate. For $(\mathfrak{x},\mathfrak{t})\in[0,N^{-1}]\times[0,N^{-2}]$, define $a_{j}^{\mathfrak{x},\mathfrak{t}}$ by
$a_{j}^{\mathfrak{x},\mathfrak{t}}=a_{j}e_{\mathbb{R}}(-j\mathfrak{x}-j^{2}\mathfrak{t}).$
Define $g^{\mathfrak{x},\mathfrak{t}}$ by
$g^{\mathfrak{x},\mathfrak{t}}(x,t)=\sum_{j\in S}a_{j}^{\mathfrak{x},\mathfrak{t}}e_{\mathbb{Q}_{p}}\Big(jx+j^{2}t\Big),$
and correspondingly set
$\mathcal{A}^{\mathfrak{x},\mathfrak{t}}=\Big\{0\leq a\leq N-1,0\leq b\leq N^{1+\sigma}-1:\big(\frac{a}{N},\frac{b}{N^{1+\sigma}}\big)\in\{x\in\mathbb{Q}_{p}^{2}:|g^{\mathfrak{x},\mathfrak{t}}(x)|>\alpha\}\Big\}.$
Note that $|a_{j}^{\mathfrak{x},\mathfrak{t}}|=|a_{j}|$ for all $\mathfrak{x},\mathfrak{t}$. In particular, (4.2) holds for all $\mathfrak{x},\mathfrak{t}$ with the same upper bound, ranging over the distinct $g^{\mathfrak{x},\mathfrak{t}}$.
We integrate:
$\begin{split}\mathcal{L}^{2}\big(U_{\alpha}\cap A(N,\sigma;\gamma)\big)&=\int_{\begin{subarray}{c}\mathfrak{x}\in[0,N^{-1}]\\ \mathfrak{t}\in[0,N^{-2}]\end{subarray}}\sum_{a=0}^{N-1}\sum_{b=0}^{N^{1+\sigma}-1}1_{U_{\alpha}}\Big(\frac{a}{N}+\mathfrak{x},\frac{b}{N^{1+\sigma}}+\mathfrak{t}\Big)\\ &=\int_{\begin{subarray}{c}\mathfrak{x}\in[0,N^{-1}]\\ \mathfrak{t}\in[0,N^{-2}]\end{subarray}}\#\mathcal{A}^{\mathfrak{x},\mathfrak{t}}\\ &=\int_{\begin{subarray}{c}\mathfrak{x}\in[0,N^{-1}]\\ \mathfrak{t}\in[0,N^{-2}]\end{subarray}}N^{-\sigma}\mu(\{x\in\mathbb{Q}_{p}^{2}:|g^{\mathfrak{x},\mathfrak{t}}(x)|>\alpha\}).\end{split}$
Applying our uniform upper bound, we conclude. ∎